OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x

OpenStudy (anonymous):

\[\frac{ x+20 }{ 2 }=3x\] \[2(\frac{ x+20 }{ 2 })=3x(2)\] This cancels out the 2 in the fraction \[x+20=6x\] -x -x 20 = 5x /5 /5 x = 4